OpenStudy (anonymous):

What is the arc length of the subtending arc for an angle of 72 degrees on a circle of radius 4?

OpenStudy (amistre64):

where is the angle located?

OpenStudy (anonymous):

\[\frac{73 \pi 8}{360}=\frac{73 \pi }{45}=5.09636 \]

OpenStudy (anonymous):

thank you guys sorry i havent been on